OpenStudy (anonymous):

Find the equillibrium solution to y'=-ty Now, I know Im supposed to just set this equal to zero, but my question is, is my answer just when y=0, or is it when y=0 OR t=0?

OpenStudy (amistre64):

y' + ty = 0

OpenStudy (amistre64):

your answer is the family of functions that make this statement true

OpenStudy (amistre64):

let y=e^(rx) and build on the product rule of derivatives to play with.

OpenStudy (amistre64):

e^(rt) but yeah ....

OpenStudy (anonymous):

yes y=0 or t=0

OpenStudy (amistre64):

if you can recall some laplace rules, that a fine method too

OpenStudy (amistre64):

or let:\[ty=\sum_0c_nt^{n+1}\] \[y'=\sum_1c_n~nt^{n-1}\] \[y'+ty=\sum_1c_n~nt^{n-1}+\sum_0c_nt^{n+1}=0\]

OpenStudy (anonymous):

@amistre64 dont we put y'= 0 to find equilibrium solution

OpenStudy (amistre64):

\[\sum_1c_n~nt^{n-1}+\sum_0c_nt^{n+1}=0\] \[\sum_1c_n~nt^{n-1}+\sum_2c_{n-2}t^{n-1}=0\] \[c_1+\sum_2c_n~nt^{n-1}+\sum_2c_{n-2}t^{n-1}=0\] \[c_1+\sum_2[c_n~n+c_{n-2}]t^{n-1}=0\] \[c_n=-\frac{c_{n-2}}{n}~:~c_1=0~;~n\ge2\] is one approach ....

OpenStudy (amistre64):

becuase that messes up the uniqueness quailfications

OpenStudy (amistre64):

no, we put: not sure what you mean by an equilibrium soltuion

OpenStudy (amistre64):

if y' = 0, then y = some constant so yeah, i spose y=0 is an equilibrium solution .... but does not represent all the solutions

OpenStudy (amistre64):

but to me it appears that what we mean by y=0 is simply that the function that y represents, is equal to 0 y(t) = 0 is a solution, but the family of solutions is y(t) = c e^(-t^2/2)

OpenStudy (anonymous):

haha am not sure i just googled and they put y' = 0 so asked

OpenStudy (amistre64):

y'=0 isnt a valid approach since y=c y' + ty = 0 0 + tc = 0, is not true for all values of c

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

We *do* put \(y'=0\), and we solve for \(y\). That gives \(0=ty~~\Rightarrow~~y=0\).

OpenStudy (anonymous):

-.-